Find the curvature, the radius of curvature and the center of curvature having a parametric equations at the given point:



r = 4cos 2theta , theta = 1/12 * pi


1
Expert's answer
2021-12-01T11:27:13-0500

r = 4cos2"\theta"


r1= "\frac{dr}{d\theta}=-8sin2\theta"

r2= "\frac{d²r}{d\theta^{2}}=-16cos2\theta"

Curvature at (r,"\theta)" = "\frac{r²+2r_{1}^{2}-rr_{2}}{(r²+r_{1}²)^{3/2}}"

r("\theta=\frac{π}{12})=4cos(\frac{π}{6})=2\sqrt{3}"

r1("\theta=\frac{π}{12})=" "-8sin(\frac{π}{6})=-4"

r2("\theta=\frac{π}{12})=" "-16cos(\frac{π}{6})=-8\sqrt{3}"

rr2 at "(\theta = \frac{π}{12})=" "-2\sqrt{3}.8\sqrt{3}=-48"

Therefore curvature at "(\theta = \frac{π}{12})="

"\frac{(2\sqrt{3})²+2(-4)²+48}{((2\sqrt{3})²+(-4)²)^{3/2}}" = "\frac{92}{(28)^{3/2}}" = "\frac{23}{2(7)^{3/2}}"

and radius of curvature = "\frac{2(7)^{3/2}}{23}"